OpenStudy (anonymous):

r = <5, 9, -4>, v = <3, 3, -8>, w = <9, -6, -1> v ⋅ w

OpenStudy (anonymous):

\[v\cdot w =3\times 9+3\times (-6)+(-8)\times (-1)\] whatever that is

OpenStudy (anonymous):

17 c: